7. Forest Components makes aircraft parts. The following transactions occurred in July. Purchased $16,910 of materials...

7.

Forest Components makes aircraft parts. The following transactions occurred in July.

  1. Purchased $16,910 of materials on account.

  2. Issued $16,720 in direct materials to the production department.

  3. Issued $1,350 of supplies from the materials inventory.

  4. Paid for the materials purchased in transaction (1) using cash.

  5. Returned $2,080 of the materials issued to production in (2) to the materials inventory.

  6. Direct labor employees earned $32,400, which was paid in cash.

  7. Purchased miscellaneous items for the manufacturing plant for $17,250 on account.

  8. Recognized depreciation on manufacturing plant of $35,200.

  9. Applied manufacturing overhead for the month.

Forest uses normal costing. It applies overhead on the basis of direct labor costs using an annual, predetermined rate. At the beginning of the year, management estimated that direct labor costs for the year would be $435,600. Estimated overhead for the year was $387,684.

The following balances appeared in the inventory accounts of Forest Components for July.

Beginning Ending
Materials Inventory ? $ 12,410
Work-in-Process Inventory ? 10,550
Finished Goods Inventory $ 2,770 6,980
Cost of Goods Sold ? 74,300

Required:

a. Prepare journal entries to record these transactions.

b. Prepare T-accounts to show the flow of costs during the period from Materials Inventory through Cost of Goods Sold.
Answer #1
predeternined overhead rate
estimated overhead/estimated direct labor costs
387684/435600
89%
TR General journal Debit Credit
1) Raw materials inventory 16,910
Accounts payable 16,910
2) Work in process inventory 16,720
Raw materials inventory 16,720
3) manufacturing overhead 1,350
Raw materials inventory 1,350
4) Accounts payable 16,910
cash 16,910
5) Raw materials inventory 2,080
work in process inventory 2,080
6) work in process inventory 32,400
cash 32,400
7) manufacturing overhead 17,250
Accounts payable 17,250
8) Manufacturing overhead 35,200
Accumulated depreciation 35,200
9) work in process inventory 31104
applied Manufacturing overhead 31,104
(32400*89%)
Materials inventory
Beg.Bal 11,490 Work in process inventory
1) 16,910 16,720 2) beg .bal 10,916
5) 2,080 1,350 3) 2) 16,720 2,080 5)
6) 32,400 78,510 Trfd to finished
9) 31104 goods
end bal 12,410
End bal 10,550
Manufacturing overhead control Applied manufacturing overhead
Beg.bal 0 Beg.Bal 0
3) 1,350 31104 9)
7) 17,250
8) 35,200
End bal 53800 end bal 31104
Accounts payable Cash
Beg.Bal 0 Beg.Bal 0
4) 16,910 16,910 1) 16,910 4)
17,250 7) 32,400 6)
End bal 17,250 End bal 49,310
Accumulated Depreciation -PP&E finished goods inventory
Beg.bal 0 Beg bal 2,770
35,200 8) goods com 78,510 74,300 Trfd to COGS
End bal 35200 end bal 6,980
cost of goods sold
Beg.Bal 0
Goods comp 74,300
End bal 74300
Answer #2

material inventory: (16,910+2080)- (16,720+1,350) = 920

12,410-920= 11,490


Work In Process: (16,720+32,400+31,104)-(2,080+78,510)=-366

366+10,550= 10,916


finished goods inventory

74,300-2,770= 71,530  +6,980=78,510